# Rydberg superatoms: An artificial quantum system for quantum information processing and quantum optics

### v1

@@ V APPLICATIOINS IN QUANTUM OPTICS
In this section, we first present a brief review of recent experimental and theoretical progress focusing on the nonlocal nonlinear responses, as induced by the dipole-dipole interactions and related to the dipole blockade effect, of cold Rydberg atoms driven into the EIT regime.Fleischhauer, Imamoglu, and Marangos (2005) Then we look back at some important proposals and schemes for achieving efficient nonlinear quantum opticsChang, Vuletić, and Lukin (2014) by considering hybrid systems with Rydberg atoms restricted in cavities, placed close to waveguides, or trapped above atomic chips.
@@
Figure 20: (a) A ladder scheme of three-level atoms that interact with a probe and a coupling field. $\mathcal{V}_{\rm vdW}$ denotes the vdW interaction between two atoms in the Rydberg state $|r\rangle$. (b) Truncated level scheme of a superatom composed of $n_{SA}$ atoms, with collectively enhanced transition amplitudes involving different numbers of probe and coupling photons. (c) Top: Probe field transmission against detuning $\Delta_{p}$ for different input probe Rabi frequencies. The thin lines are experimental curves from Ref. Pritchard et al., 2010 while thick lines are numerical simulations based on the superatom model. Bottom: Corresponding intensity correlation function of the transmitted probe field at the end of the sample. Reprinted with permission from Petrosyan et al., Phys. Rev. Lett. 107(21), 213601 (2011). Copyright 2011 American Physical Society.Petrosyan, Otterbach, and Fleischhauer (2011)
Figure 21: (a) Rydberg-blockade-mediated interaction between slow photons. (b) Transmission versus probe detuning at various incoming photon rates. (c) Data points show photon–photon correlation function $g^{(2)}(\tau)$ at EIT resonance. Reprinted with permission from Peyronel et al., Nature 488(7409), 57-60 (2012). Copyright 2012 Springer Nature.Peyronel et al. (2012)
V.1 Nonlocal nonlinear optics
Typical EIT experiments are performed with atoms in low-lying states driven into the three-level $\Lambda$ configuration involving an intermediate short-lived excited state and two long-lived ground states. When a ground state is replaced with a Rydberg state, it is viable to realize an intriguing transfer from the linear EIT in the $\Lambda$ configuration to the nonlinear EIT in the ladder configuration, as a result of both long-lifetime and strong-interaction features of highly excited Rydberg atoms.
In 2010, Adams’ group observed in ${}^{87}$b atoms a cooperative optical nonlinearity arising from Rydberg interactions and manifested itself as a gradual reduction of the probe transmission in the EIT window when the probe intensity increasesPritchard et al. (2010) as shown in Fig. 19. This is studied considering a ladder EIT system involving ground $|g\rangle$, intermediate $|e\rangle$, and upper $|r\rangle$ levels and is explained by a simple two-atom model where one atom contributes to the three-level EIT system while the other acts as a two-level absorbing system due to dipole blockade. This fascinating experiment soon is better explained by a mean-field superatom model where the collective states of $n_{SA}=\frac{4}{3}\pi\rho R_{SA}^{3}$ atoms within a blockade sphere of radius $R_{SA}=(C_{6}\gamma_{e}/\Omega_{c}^{2})^{1/6}$ are considered but appropriately truncatedPetrosyan, Otterbach, and Fleischhauer (2011) as shown in Figs. 20(a) and 20(b). This truncation is done based on two facts: i.e., (i) at most one atom is allowed to be in the upper Rydberg state and (ii) the excitation to the intermediate state $|e\rangle$ is well suppressed due to quantum destructive interference. Then it is appropriate to consider only the three lowest-order collective states $|G\rangle=|g_{1},g_{2},...,g_{n_{SA}}\rangle$, $|E^{(1)}\rangle=\frac{1}{\sqrt{n_{SA}}}\sum_{j}^{n_{SA}}|g_{1},...,e_{j},...g_{n_{SA}}\rangle$, and $|R^{(1)}\rangle=\frac{1}{\sqrt{n_{SA}}}\sum_{j}^{n_{SA}}|g_{1},...,r_{j},...g_{n_{SA}}\rangle$ for each superatom as one attempts to estimate the Rydberg excitation probability $\Sigma_{RR}$. With $\Sigma_{RR}$, it is easy to calculate the conditional polarizability $\alpha=\Sigma_{RR}\alpha_{TLA}+(1-\Sigma_{RR})\alpha_{EIT}$ with $\alpha_{TLA}$ and $\alpha_{EIT}$ being the polarizabilities for the two-level absorbing and ladder EIT systems, jointly determining the probe field propagation in the Rydberg medium. Such a computationally efficient model is in quantitative agreement with the experimental results and also predicts in Fig. 20(c) a remarkable antibunching effect in terms of the intensity correlation function $g^{(2)}_{p}(L)$ observed at the sample exit, accompanied by the transmission reduction due to cooperative optical nonlinearity.
Soon afterward, higher-order collective states are included in an improved superatom model by going beyond the weak-probe approximation, which just results in quantitative differences, especially as the probe and coupling fields are off two-photon resonance, yet without changing main conclusions.Liu et al. (2014) Such a superatom model has been further extended to realize the polarization-selective nonlinearity and cooperative nonlinear grating unattainable in normal atoms.Wu, Artoni, and La Rocca (2015); Liu et al. (2016); Asghar et al. (2016) A semianalytical rate-equation model has also been developed for revealing a universal relation between the nonlinear optical susceptibility and the Rydberg excitation density as interatomic coherences are well suppressed due to strong dephasing effects or very different Rabi frequencies of the probe and coupling fields.Gärttner et al. (2014) Moreover, we note that a many-body iterative approach has been developed to account for the interparticle Rydberg correlations, though it is applicable only for dilute samples with less than one atom inside a superatom so that Rydberg interactions between different superatoms become dominant to reduce the many-body computational complexity. This alternative approach allows one to separate a local linear susceptibility and a nonlocal nonlinear susceptibility comparable in magnitude, hence achieving huge nonlocal optical nonlinearities beneficial, e.g., to largely enhance the Goos-Hanchen lateral shift and correlated biphoton generation.Bai and Huang (2016); Bai, Li, and Huang (2019); Zheng et al. (2022); Zhao et al. (2023)
Figure 22: (a) A light storage and retrieval experiment based on the nonlinear Rydberg EIT technique. (b) Normalized and background subtracted counts of an input probe pulse when slowly propagating or stored for $t_{s}=600$ ns. (c) Output photon number normalized by linear process efficiency $T$ against input photon number for the slow-light case and two storage times. Straight lines represent the linear behavior (oblique) and the saturation level (horizontal). Reprinted with permission from Distante et al., Phys. Rev. Lett. 117(11), 113001 (2016). Copyright 2016 American Physical Society.Distante et al. (2016)
An important application of the linear EIT in normal atoms is to realize the dynamically reversible storage and retrieval of a (classical or quantum) probe pulse by adiabatically switching off and on a control field. Such an important coherent control of the slow-light propagation and storage dynamics has been extended to Rydberg atoms exhibiting nonlocal nonlinear EIT responses. Many experiments show that storing an input probe pulse as a collective Rydberg excitation strongly enhances the Rydberg-mediated interaction compared to that in the slow-light dark-state polariton (DSP) case. For example, Distante et al. observed a strong enhancement of interatomic interactions due to the largely reduced polariton group velocity for short storage times while a weak enhancement dominated by Rydberg-induced dephasing of the polariton’s multiparticle components for longer storage times.Distante et al. (2016) This conclusion is supported by Fig. 22 where the output photon number $N_{\rm out}$ depends on the input photon number $N_{\rm in}$ clearly in a nonlinear way and becomes saturated for a sufficiently large $N_{\rm in}$ as described by the input-output relation $N_{\rm out}=N_{\rm max}T(1-e^{-N_{\rm in}/N_{\rm max}})$ with $T$ being the linear process efficiency for small $N_{\rm in}$. This nonlinear light storage experiment can be well explained by extending the superatom model from its steady-state solutions to the dynamic evolution case Tian et al. (2018) which further predicts an inhomogeneous nonclassical antibunching feature of the retrieved light signal. Another effective method has also been developed to reveal the correlated photon dynamics in dissipative Rydberg media valid in the limit of a large optical depth per blockade radius.Gorshkov, Nath, and Pohl (2013); Zeuthen et al. (2017)
Figure 23: Experimental realization of contactless nonlinear optics mediated by long-range Rydberg interactions. Reprinted with permission from Busche et al., Nature Physics 13(7), 655-658 (2017). Copyright 2017 Springer Nature.Busche et al. (2017)
Even in the absence of a common medium, creating effective photon-photon interactions remains feasible.Yang et al. (2016) Busche et al.’s study delved into a novel approach to such photon interactions utilizing Rydberg interactions.Busche et al. (2017) As depicted in Fig. 23, photons stored as collective Rydberg excitations in spatially separated optical media interact through van der Waals interactions, inducing non-uniform phase shifts that alter the retrieval modes of the photons. These interaction-induced phase shifts are pivotal in shaping the overall interaction between the photons, resulting in observable effects like reduced retrieval in the photons’ initial modes and spatial anti-correlations in observed photon statistics. The experimental arrangement utilized highly focused signal beams and tiny optical tweezers to confine cold atoms, demonstrating the high level of precision and control in investigating non-contact nonlinear optics. Results exhibit strong agreement between theory and experiment across varying interaction strengths and distances, indicating the feasibility of implementing contactless effective photon-photon interactions in practical photonic devices. This contactless characteristic of interactions unveils new opportunities for scalable multichannel photonic devices, exploration of strongly correlated many-body dynamics using light, and the development of advanced quantum technologies based on photon interactions mediated by Rydberg states.
V.2 Single-photon engineering
Rigid dipole blockade effect in Rydberg atoms allows one to attain efficient single-photon engineering, including single-photon generation and storage due to collectively enhanced strong couplings between single Rydberg superatoms and few-photon light pulses.Paris-Mandoki et al. (2017) In Ref. Saffman and Walker, 2002, Saffman and Walker conducted a detailed analysis of collective emission and proposed the protocol for realizing deterministic single-photon source using Rydberg superatom. Instead of using Rydberg EIT to store a single photon, they show that the Rydberg superatom can be coherently prepared by using two-photon excitation to drive a many-body Rabi oscillation, hence efficiently transferred into a highly directional single photon. Honer et al. further examined photon interactions with a Rydberg superatom under controlled dephasing conditions and discover that induced dephasing may dictate the absorption of a single photon from any probing field,Honer et al. (2011) which was recently employed in the controlled multi-photon subtraction process.Stiesdal et al. (2021)
Figure 24: Single-photon source with Rydberg superatom. (a) Illustration of experimental protocol. (b) Relevant atomic levels. (c) Measured second-order-correlation function $g^{2}(0)$ as a function of the effective principal quantum number. Reprinted with permission from Dudin et al., Science 336(6083), 887–889 (2012). Copyright 2012 The American Association for the Advancement of Science.Dudin and Kuzmich (2012)
A deterministic single-photon source based on Rydberg superatom was experimentally realized in 2012. Dudin and Kuzmich (2012) As shown in Fig. 24, Dudin and Kuzmich performed 780 nm - 480 nm two-photon Rydberg excitation and photon retrieval in an ensemble of ${}^{87}$b atoms confined within the blockade radius. When high-lying Rydberg states ($n\sim 100$) are used, the measured second-order-correlation function $g^{2}(0)$ of the retrieved photon approaches zero, indicating the successful preparation of a single photon. Further research in Rydberg single-photon source focused on analyzing and improving the quality of generated single photons.Lampen et al. (2019); Ornelas-Huerta et al. (2020); Padrón-Brito et al. (2021); Shi et al. (2022) The state-of-the-art Rydberg single-photon sources feature near-unity purity and indistinguishability, which are comparable to those in other leading single-photon physical platformsAharonovich, Englund, and Toth (2016) and have been used to achieve a high-fidelity photonic quantum logic gate using linear optical approach.Shi et al. (2022)
Figure 25: Transition $|u\rangle\rightarrow|d\rangle$ of a source atom (top left) is coupled nonresonantly to transition $|i\rangle\rightarrow|s\rangle$ of the medium atoms (top right) via DDE interaction $D$. A laser pulse $\Omega$ couples ground state $|g\rangle$ to Rydberg state $|i\rangle$ and leads to the transition $|g\rangle\rightarrow|s\rangle$ with the help of DDE transition $|i\rangle|u\rangle\rightarrow|s\rangle|d\rangle$. The resulting single Rydberg excitation exhibits the profile $S(\textbf{r})\propto D(\textbf{r}-\textbf{r}_{s})$ and can be converted into a propagating photon $\mathcal{E}$ by applying the control field $\Omega_{c}$. Reprinted with permission from Petrosyan et al., Phys. Rev. Lett. 121(12), 123605 (2018). Copyright 2018 American Physical Society.Petrosyan and Mølmer (2018)
Figure 26: (a) Level scheme, (b) simplified schematic, and (c) pulse sequence for the realization of an all-optical transistor with $U(r)$ denoting the dipole-dipole interaction between two atoms of distance $r$ and in states $|r_{g}\rangle$ and $|r_{s}\rangle$, respectively. (d) Absorption spectrum of the source field (dots) exhibiting an EIT window at $\delta_{s}=0$ and that of the gate field (circles) around $\delta_{g}=40$ MHz. The contrast of the switches (red solid) against mean numbers of (e) incident and (f) stored gate photons in coherent states is presented with fundamental limits (pink dashed) set by the corresponding photon statistics. (g) Gate photon optical gain and (h) source photon transfer function with $\overline{N}_{\rm g,in}=0$ (red) and $\overline{N}_{\rm g,in}=0.75$ (blue) for coherent states, while $N_{\rm in,Fock}=1$ (green) and $N_{\rm st,Fock}=1$ (black) for Fock states. Reprinted with permission from Gorniaczyk et al., Phys. Rev. Lett. 113(5), 053601 (2014). Copyright 2014 American Physical Society.Gorniaczyk et al. (2014)
The dipole blockade and exchange mechanisms described above used for generating single-photon sources can also be extended to design usually unattainable single-photon devices, including single-photon switch,Baur et al. (2014) transistor, Gorniaczyk et al. (2014); Tiarks et al. (2014) and absorber. Paris-Mandoki et al. (2017) For instance, one may store a gate light pulse containing only one photon on average as a collective Rydberg excitation in cold atoms, in the presence of which the transmission of the subsequent target pulse is largely suppressed due to dipole blockade.Baur et al. (2014) Such a single-photon switch offers many interesting prospects ranging from quantum communication to quantum computing, such as heralding a successful quantum memory event, nondestructively detecting an optical photon, and building a photonic quantum-logic gate. Moreover, it is viable to realize an all-optical single-photon transistor by mapping gate and source (target) photons to strongly interacting Rydberg excitations with different principal quantum numbers Gorniaczyk et al. (2014) or by suppressing and enhancing the transmission of target (source) photons through dipole blockade and Förster resonances, respectively.Tiarks et al. (2014)
To be more specific, we can see from Figs. 26(a)-26(d) that an all-optical transistor can be achieved with the following two procedures. First, store photons in a gate field $\mathcal{E}_{g}$ as collective Rydberg excitation in a sample of cold ${}^{87}$b atoms by coupling the ground state $|g\rangle$ with the Rydberg state $|r_{g}\rangle$ through a control field $\Omega_{g}$. Second, a source field $\mathcal{E}_{s}$ is sent through the same medium at a reduced group velocity due to the resonant EIT provided by the control field $\Omega_{s}$ coupling to the Rydberg state $|r_{s}\rangle$. Then individual source photons can travel with little absorption through the atomic sample in the absence of a gate excitation for a low source photon rate or are strongly attenuated by the scattering from the intermediate state $|e\rangle$ in the blockade regime where Rydberg interactions between states $|r_{g}\rangle$ and $|r_{s}\rangle$ result in a two-level absorbing system. This realizes a transistor working at the single-photon level with its main performance illustrated in Figs. 26(e)-26(h). First, we can see from Fig. 26(e) that the switch contrast $C$ plotted against the mean number of incident gate photons $\overline{N}_{g,{\rm in}}$ is clearly below the fundamental limit set by the coherent-state photon statistics. Fig. 26(f) shows, however, that $C$ plotted against $\overline{N}_{g,{\rm st}}$ (the mean number of stored gate photons) gets much closer to the corresponding fundamental limit. It is more important to note that, on average, about $3$ input or $2$ stored gate photons are already enough to achieve a switch contrast over $0.8$. Fig. 26(g) further shows that the optical gain is up to $20\sim 30$ (limited by the self-blocking effect of the source photons) where a coherent input with $\overline{N}_{g,{\rm in}}=0.75(3)$ gate photons can be used to effectively control the transmission of $30\sim 50$ source photons. Finally, we find in Fig. 26(h) a nonlinear saturation behavior due to self-blockade in terms of the source photon transfer function from $\overline{N}_{s,{\rm in}}$ to $\overline{N}_{s,{\rm out}}$. Hence a much higher optical gain is expected to occur provided the source photon loss due to self-blockade can be overcome.
V.3 Hybrid Rydberg system

### v2

@@ V APPLICATIOINS IN QUANTUM OPTICS
In this section, we first present a brief review of recent experimental and theoretical progress focusing on the nonlocal nonlinear responses, as induced by the dipole-dipole interactions and related to the dipole blockade effect, of cold Rydberg atoms driven into the EIT regime.Fleischhauer, Imamoglu, and Marangos (2005); Finkelstein et al. (2023) Then we look back at some important proposals and schemes for achieving efficient nonlinear quantum opticsChang, Vuletić, and Lukin (2014) by considering hybrid systems with Rydberg atoms restricted in cavities, placed close to waveguides, or trapped above atomic chips.
Figure 19: (a) Schematic of a Rydberg EIT experiment with cold 87Rb atoms driven by a probe field on the lower $|g\rangle\leftrightarrow|e\rangle$ transition and a coupling field on the upper $|e\rangle\leftrightarrow|r\rangle$ transition in the ladder configuration. (b) Nonlinear EIT spectra of the probe field where the transmission peak on resonance ($\Delta_{p}=0$) reduces as probe Rabi frequency $\Omega_{p}$ increases. The experiment is done for the $|r\rangle=|60S_{1/2}\rangle$ state at atomic density $\rho=1.2\pm 0.1\times 10^{10}$ cm-3. Reproduced with permission from Pritchard et al., Phys. Rev. Lett. 105(19), 193603 (2010). Copyright 2010 American Physical Society.Pritchard et al. (2010)
Figure 20: (a) A ladder scheme of three-level atoms that interact with a probe and a coupling field. $\mathcal{V}_{\rm vdW}$ denotes the vdW interaction between two atoms in the Rydberg state $|r\rangle$. (b) Truncated level scheme of a superatom composed of $n_{SA}$ atoms, with collectively enhanced transition amplitudes involving different numbers of probe and coupling photons. (c) Top: Probe field transmission against detuning $\Delta_{p}$ for different input probe Rabi frequencies. The thin lines are experimental curves from Ref. Pritchard et al., 2010 while thick lines are numerical simulations based on the superatom model. Bottom: Corresponding intensity correlation function of the transmitted probe field at the exit from the medium. Reproduced with permission from Petrosyan et al., Phys. Rev. Lett. 107(21), 213601 (2011). Copyright 2011 American Physical Society.Petrosyan, Otterbach, and Fleischhauer (2011)
Figure 21: (a) Rydberg-blockade-mediated interaction between slow photons. (b) Transmission versus probe detuning at various incoming photon rates. (c) Data points show photon–photon correlation function $g^{(2)}(\tau)$ at EIT resonance. Reproduced with permission from Peyronel et al., Nature 488(7409), 57-60 (2012). Copyright 2012 Springer Nature.Peyronel et al. (2012)
V.1 Nonlocal nonlinear optics
Typical EIT experiments are performed with atoms in low-lying states driven into the three-level $\Lambda$ configuration involving an intermediate short-lived excited state and two long-lived ground states.Harris, Field, and Imamoğlu (1990); Boller, Imamoğlu, and Harris (1991) When a ground state is replaced with a Rydberg state, it is viable to realize an intriguing transfer from the linear EIT in the $\Lambda$ configuration to the nonlinear EIT in the ladder configuration, as a result of both long-lifetime and strong-interaction features of highly excited Rydberg atoms.
In 2010, Adams’ group observed in 87Rb atoms a cooperative optical nonlinearity arising from Rydberg interactions and manifesting itself as a gradual reduction of the probe transmission in the EIT window when the probe intensity increasesPritchard et al. (2010) as shown in Fig. 19. The study was performed with an EIT ladder system involving ground $|g\rangle$, intermediate $|e\rangle$, and upper $|r\rangle$ levels and was explained by a simple two-atom model where one atom contributes to the three-level EIT system while the other acts as a two-level absorbing system due to dipole blockade. This fascinating experiment soon was better explained by a mean-field superatom model where the collective states of $n_{SA}=\frac{4}{3}\pi\rho R_{SA}^{3}$ atoms within a blockade sphere of radius $R_{SA}=(C_{6}\gamma_{e}/\Omega_{c}^{2})^{1/6}$ defined as the interatomic distance at which the vdWs interaction equals the half-width of steady-state population of a single atomic Rydberg state,Qian et al. (2013) are considered but appropriately truncatedPetrosyan, Otterbach, and Fleischhauer (2011) as shown in Figs. 20(a) and 20(b). This truncation is done based on two facts: i.e., (i) at most one atom is allowed to be in the upper Rydberg state and (ii) the excitation to the intermediate state $|e\rangle$ is well suppressed due to quantum destructive interference. Then it is appropriate to consider only the three lowest-order collective states $|G\rangle=|g_{1},g_{2},...,g_{n_{SA}}\rangle$, $|E^{(1)}\rangle=\frac{1}{\sqrt{n_{SA}}}\sum_{j}^{n_{SA}}|g_{1},...,e_{j},...g_{n_{SA}}\rangle$, and $|R^{(1)}\rangle=\frac{1}{\sqrt{n_{SA}}}\sum_{j}^{n_{SA}}|g_{1},...,r_{j},...g_{n_{SA}}\rangle$ for each superatom as one attempts to estimate the Rydberg excitation probability $\Sigma_{RR}$. With $\Sigma_{RR}$, it is easy to calculate the conditional polarizability $\alpha=\Sigma_{RR}\alpha_{TLA}+(1-\Sigma_{RR})\alpha_{EIT}$ with $\alpha_{TLA}$ and $\alpha_{EIT}$ being the polarizabilities for the two-level absorbing and ladder EIT systems, jointly determining the probe field propagation in the Rydberg medium. Such a computationally efficient model is in quantitative agreement with the experimental results and also predicts in Fig. 20(c) a remarkable antibunching effect in terms of the intensity correlation function $g^{(2)}_{p}(L)$ obtained at the exit from the medium, accompanied by a transmission reduction due to cooperative optical nonlinearity.
Soon afterward, higher-order collective states were included in an improved superatom model by going beyond the weak-probe approximation, which just results in quantitative differences, especially as the probe and coupling fields are off two-photon resonance, yet without changing main conclusions.Liu et al. (2014) Such a superatom model has been further extended to realize the polarization-selective nonlinearity and cooperative nonlinear grating unattainable in normal atoms.Wu, Artoni, and La Rocca (2015); Liu et al. (2016); Asghar et al. (2016) A semianalytical rate-equation model has also been developed for revealing a universal relation between the nonlinear optical susceptibility and the Rydberg excitation density as interatomic coherence is well suppressed due to strong dephasing effects or very different Rabi frequencies of the probe and coupling fields.Gärttner et al. (2014) Moreover, we note that a many-body iterative approach has been developed to account for the interparticle Rydberg correlations, though it is applicable only for dilute samples with less than one atom inside a superatom so that Rydberg interactions between different superatoms become dominant to reduce the many-body computational complexity. This alternative approach allows one to separate a local linear susceptibility and a nonlocal nonlinear susceptibility comparable in magnitude, hence achieving huge nonlocal optical nonlinearities beneficial, e.g., to largely enhance the Goos-Hanchen lateral shift and correlated biphoton generation.Bai and Huang (2016); Bai, Li, and Huang (2019); Zheng et al. (2022); Zhao et al. (2023)
Figure 22: (a) A light storage and retrieval experiment based on the nonlinear Rydberg EIT technique. (b) Normalized and background subtracted counts of an input probe pulse when slowly propagating or stored for $t_{s}=600$ ns. (c) Output photon number normalized by linear process efficiency $T$ against input photon number for the slow-light case and two storage times. Straight lines represent the linear behavior (oblique) and the saturation level (horizontal). Reproduced with permission from Distante et al., Phys. Rev. Lett. 117(11), 113001 (2016). Copyright 2016 American Physical Society.Distante et al. (2016)
An important application of the linear EIT in normal atoms is to realize the dynamically reversible storage and retrieval of a (classical or quantum) probe pulse by adiabatically switching off and on a control field. Such an important coherent control of the slow-light propagation and storage dynamics has been extended to Rydberg atoms exhibiting nonlocal nonlinear EIT responses. Many experiments show that storing an input probe pulse as a collective Rydberg excitation strongly enhances the Rydberg-mediated interaction compared to that in the slow-light dark-state polariton case. For example, Distante et al. observed a strong enhancement of interatomic interactions due to the largely reduced polariton group velocity for short storage times while a weak enhancement dominated by Rydberg-induced dephasing of the polariton’s multiparticle components for longer storage times.Distante et al. (2016) This conclusion is supported by Fig. 22 where the output photon number $N_{\rm out}$ depends on the input photon number $N_{\rm in}$ clearly in a nonlinear way and becomes saturated for a sufficiently large $N_{\rm in}$ as described by the input-output relation $N_{\rm out}=N_{\rm max}T(1-e^{-N_{\rm in}/N_{\rm max}})$ with $T$ being the linear process efficiency for small $N_{\rm in}$. This nonlinear light storage experiment can be well explained by extending the superatom model from its steady-state solutions to the dynamic evolution case Tian et al. (2018) which further predicts an inhomogeneous nonclassical antibunching feature of the retrieved light signal. Another effective method has also been developed to reveal the correlated photon dynamics in dissipative Rydberg media valid in the limit of a large optical depth per blockade radius.Gorshkov, Nath, and Pohl (2013); Zeuthen et al. (2017)
Figure 23: Experimental realization of contactless nonlinear optics mediated by long-range Rydberg interactions. Reproduced with permission from Busche et al., Nature Physics 13(7), 655-658 (2017). Copyright 2017 Springer Nature.Busche et al. (2017)
Even in the absence of a common medium, creating effective photon-photon interactions remains feasible.Yang et al. (2016) Busche et al.’s study delved into a novel approach to such photon interactions utilizing Rydberg interactions.Busche et al. (2017) As depicted in Fig. 23, photons stored as collective Rydberg excitations in spatially separated optical media interact through vdw interactions, inducing non-uniform phase shifts that alter the retrieval modes of the photons. These interaction-induced phase shifts are pivotal in shaping the overall interaction between the photons, resulting in observable effects like reduced retrieval in the photons’ initial modes and spatial anti-correlations in observed photon statistics. The experimental arrangement utilized highly focused signal beams and tiny optical tweezers to confine cold atoms, demonstrating the high level of precision and control in investigating non-contact nonlinear optics. Results exhibit strong agreement between theory and experiment across varying interaction strengths and distances, indicating the feasibility of implementing contactless effective photon-photon interactions in practical photonic devices. This contactless characteristic of interactions unveils new opportunities for scalable multichannel photonic devices, exploration of strongly correlated many-body dynamics using light, and the development of advanced quantum technologies based on photon interactions mediated by Rydberg states.
V.2 Single-photon engineering
Rigid dipole blockade effect in Rydberg atoms allows one to attain efficient single-photon engineering, including single-photon generation and storage due to collectively enhanced strong couplings between single Rydberg superatoms and few-photon light pulses.Paris-Mandoki et al. (2017) In Ref. Saffman and Walker, 2002, Saffman and Walker conducted a detailed analysis of collective emission and proposed the protocol for realizing deterministic single-photon source using Rydberg superatom. Instead of using Rydberg EIT to store a single photon, they show that the Rydberg superatom can be coherently prepared by using two-photon excitation to drive a many-body Rabi oscillation, hence efficiently transferred into a highly directional single photon. Honer et al.Honer et al. (2011) and Tresp et al.Tresp et al. (2016) further examined photon interactions with a Rydberg superatom under controlled dephasing conditions and discovered that induced dephasing may dictate the absorption of a single photon from any probing field, which was recently employed in the controlled multi-photon subtraction process.Stiesdal et al. (2021)
Figure 24: Single-photon source with Rydberg superatom. (a) Illustration of experimental protocol. (b) Relevant atomic levels. (c) Measured second-order-correlation function $g^{2}(0)$ as a function of the effective principal quantum number. Reproduced with permission from Dudin et al., Science 336(6083), 887–889 (2012). Copyright 2012 The American Association for the Advancement of Science.Dudin and Kuzmich (2012)
A deterministic single-photon source based on Rydberg superatom was experimentally realized in 2012. Dudin and Kuzmich (2012) As shown in Fig. 24, Dudin and Kuzmich performed 780 nm-480 nm two-photon Rydberg excitation and photon retrieval in an ensemble of 87Rb atoms confined within the blockade radius. When high-lying Rydberg states ($n\sim 100$) are used, the measured second-order-correlation function $g^{2}(0)$ of the retrieved photon approaches zero, indicating the successful preparation of a single photon. Further research in Rydberg single-photon source focused on analyzing and improving the quality of generated single photons.Lampen et al. (2019); Ornelas-Huerta et al. (2020); Padrón-Brito et al. (2021); Shi et al. (2022) The state-of-the-art Rydberg single-photon sources feature near-unity purity and indistinguishability, which are comparable to those in other leading single-photon physical platformsAharonovich, Englund, and Toth (2016) and have been used to achieve a high-fidelity photonic quantum logic gate using linear optical approach.Shi et al. (2022)
Figure 25: Transition $|u\rangle\rightarrow|d\rangle$ of a source atom (top left) is coupled nonresonantly to transition $|i\rangle\rightarrow|s\rangle$ of the medium atoms (top right) via DDE interaction $D$. A laser pulse $\Omega$ couples ground state $|g\rangle$ to Rydberg state $|i\rangle$ and leads to the transition $|g\rangle\rightarrow|s\rangle$ with the help of DDE transition $|i\rangle|u\rangle\rightarrow|s\rangle|d\rangle$. The resulting single Rydberg excitation exhibits the profile $S(\textbf{r})\propto D(\textbf{r}-\textbf{r}_{s})$ and can be converted into a propagating photon $\mathcal{E}$ by applying the control field $\Omega_{c}$. Reproduced with permission from Petrosyan et al., Phys. Rev. Lett. 121(12), 123605 (2018). Copyright 2018 American Physical Society.Petrosyan and Mølmer (2018)
Figure 26: (a) Level scheme, (b) simplified schematic, and (c) pulse sequence for the realization of an all-optical transistor with $U(r)$ denoting the dipole-dipole interaction between two atoms of distance $r$ and in states $|r_{g}\rangle$ and $|r_{s}\rangle$, respectively. (d) Absorption spectrum of the source field (dots) exhibiting an EIT window at $\delta_{s}=0$ and that of the gate field (circles) around $\delta_{g}=40$ MHz. The contrast of the switches (red solid) against mean numbers of (e) incident and (f) stored gate photons in coherent states is presented with fundamental limits (pink dashed) set by the corresponding photon statistics. (g) Gate photon optical gain and (h) source photon transfer function with $\overline{N}_{\rm g,in}=0$ (red) and $\overline{N}_{\rm g,in}=0.75$ (blue) for coherent states, while $N_{\rm in,Fock}=1$ (green) and $N_{\rm st,Fock}=1$ (black) for Fock states. Reproduced with permission from Gorniaczyk et al., Phys. Rev. Lett. 113(5), 053601 (2014). Copyright 2014 American Physical Society.Gorniaczyk et al. (2014)
The dipole blockade and exchange mechanisms described above used for generating single-photon sources can also be extended to design usually unattainable single-photon devices, including single-photon switch,Baur et al. (2014) single-photon transistor, Gorniaczyk et al. (2014); Tiarks et al. (2014) and single-photon absorber. Tresp et al. (2016); Honer et al. (2011) For instance, one may store a gate light pulse containing only one photon on average as a collective Rydberg excitation in cold atoms, in the presence of which the transmission of the subsequent target pulse is largely suppressed due to dipole blockade.Baur et al. (2014) Such a single-photon switch offers many interesting prospects ranging from quantum communication to quantum computing, such as heralding a successful quantum memory event, nondestructively detecting an optical photon, and building a photonic quantum-logic gate. Moreover, it is viable to realize an all-optical single-photon transistor by mapping gate and source (target) photons to strongly interacting Rydberg excitations with different principal quantum numbers Gorniaczyk et al. (2014) or by suppressing and enhancing the transmission of target (source) photons through dipole blockade and Förster resonances, respectively.Tiarks et al. (2014)
To be more specific, we can see from Figs. 26(a)-26(d) that an all-optical transistor can be achieved with the following two procedures. First, store photons in a gate field $\mathcal{E}_{g}$ as collective Rydberg excitation in a sample of cold 87Rb atoms by coupling the ground state $|g\rangle$ with the Rydberg state $|r_{g}\rangle$ through a control field $\Omega_{g}$. Second, a source field $\mathcal{E}_{s}$ is sent through the same medium at a reduced group velocity due to the resonant EIT provided by the control field $\Omega_{s}$ coupling to the Rydberg state $|r_{s}\rangle$. Then individual source photons can travel with little absorption through the atomic sample in the absence of a gate excitation for a low source photon rate or are strongly attenuated by the scattering from the intermediate state $|e\rangle$ in the blockade regime where Rydberg interactions between states $|r_{g}\rangle$ and $|r_{s}\rangle$ result in a two-level absorbing system. This realizes a transistor working at the single-photon level with its main performance illustrated in Figs. 26(e)-26(h). First, we can see from Fig. 26(e) that the switch contrast $C$ plotted against the mean number of incident gate photons $\overline{N}_{g,{\rm in}}$ is clearly below the fundamental limit set by the coherent-state photon statistics. Figure 26(f) shows, however, that $C$ plotted against $\overline{N}_{g,{\rm st}}$ (the mean number of stored gate photons) gets much closer to the corresponding fundamental limit. It is more important to note that, on average, about $3$ input or $2$ stored gate photons are already enough to achieve a switch contrast over $0.8$. Figure 26(g) further shows that the optical gain is up to $20\sim 30$ (limited by the self-blocking effect of the source photons) where a coherent input with $\overline{N}_{g,{\rm in}}=0.75(3)$ gate photons can be used to effectively control the transmission of $30\sim 50$ source photons. Finally, we find in Fig. 26(h) a nonlinear saturation behavior due to self-blockade in terms of the source photon transfer function from $\overline{N}_{s,{\rm in}}$ to $\overline{N}_{s,{\rm out}}$. Hence a much higher optical gain is expected to occur provided the source photon loss due to self-blockade can be overcome.
@@ V.2 Single-photon engineering
V.3 Hybrid Rydberg systems